OpenStudy (lilsis76):

Determine all of the solutions in the interval 0< Theta < 360. cos(Q/2) = 1/2 Q is theta

OpenStudy (lilsis76):

@zepdrix

OpenStudy (anonymous):

120,360

OpenStudy (anikay):

oh my, beaten to the punch gj

OpenStudy (lilsis76):

haha okay so then hmmm

OpenStudy (lilsis76):

how would that be? @zepdrix i dont get it. i see cos Q/2 that would mean cos = 1/2 right? and if its positive it would be in the 1st and 4th Quadrant?

OpenStudy (anikay):

well, now that I think about it, I think raja might be wrong

OpenStudy (anikay):

it should have 4 solutions right?

OpenStudy (lilsis76):

im not sure, it just says to determine all solutions in the interval 0-->360

OpenStudy (anonymous):

yes!

zepdrix (zepdrix):

Using Q? Oooo I like that :) I'm always having trouble finding a suitable letter for theta in text, that works quite well XD

OpenStudy (anikay):

^concur

OpenStudy (lilsis76):

lol ya, i got tired ofputting theata.

OpenStudy (anonymous):

30

OpenStudy (lilsis76):

okay is it 60 and 300 degrees?

OpenStudy (anonymous):

no tetha is 30

zepdrix (zepdrix):

\[\large \cos(\theta/2)=\frac{1}{2}\]We need to solve for theta, which involves a little more than just finding a special angle. If it helps, think of the ENTIRE inside of cosine as an angle, like this,\[\large \cos(\phi)=\frac{1}{2}\]Where is phi =1/2? Which special angles?

OpenStudy (anikay):

look, as I solve it, it's \[\theta=2\cos^{-1} (1/2)\]

OpenStudy (anikay):

the thing is, that only gives you 120. but it's a good starting point

OpenStudy (lilsis76):

mmmmm... okay, well first. special angles is like 45-45-90, and 30-60-90, right?

zepdrix (zepdrix):

The special angles are 0,30, 45, 60, 90 In radians 0, pi/6, pi/4, pi/3, pi/2 and so on... Remember those? :) We're not talking about special triangles :D Special angles on the unit circle.

OpenStudy (lilsis76):

oh....okay i gotcha now on special angles. lol im totaly flipped around in this class

OpenStudy (lilsis76):

okay then so..cosQ/2 = 1/2 right? so cos would be a positive

OpenStudy (anikay):

I think 120 is it.....

zepdrix (zepdrix):

yes the cosine is producing a positive value, so we're concerned with angles in the 1st and 4th quadrant right? :O Cuz that's where cosine is positive.

OpenStudy (lilsis76):

yes :) 1st and 4th

OpenStudy (anikay):

so, if we find the reference angle, then what would the two angles be?

OpenStudy (lilsis76):

mmm....pi/3 and 5pi/6?

OpenStudy (anikay):

yes to the first, the last, not so much, that's in the second quadrant

OpenStudy (anikay):

or am I failing?

zepdrix (zepdrix):

lol :) They should both either be /6 or /3 :D you should get similar angles in that sense. Also, the domain of theta was given to us in DEGREES. let's try to keep it in degrees :)

OpenStudy (lilsis76):

oh. haha i mean 5pi/3

OpenStudy (anikay):

With zepdrix, but you are right

OpenStudy (lilsis76):

so the second answer would be 5pi/3?

OpenStudy (anikay):

but, the problem is, we need that angle to be doubled to match our equation of theta/2

OpenStudy (lilsis76):

hmmm...ok okkay, so doubled from pi/3, so pi/3 and 5pi/6?

OpenStudy (anikay):

uh, lets throw this back to degrees shall we? pi/3 = 60 and 5pi/3 = 300

OpenStudy (lilsis76):

DANG! haha okay, so first we have 60 and u said to double it, so that would be 120, 120 would be.....that 5pi/6

OpenStudy (anikay):

no, 5pi/6 = 150 degrees

OpenStudy (lilsis76):

haha shoot what am i doing wrong

OpenStudy (anikay):

here's a tip. every pi/6 = 30 degrees

OpenStudy (lilsis76):

pi/6 pi/4 pi/3 pi/2 4pi/6div by 2 = 2pi/3 OH!.....haha okay so pi/3 and 2 pi/3???

OpenStudy (anikay):

just plug those into your original equation and see what you get :P

OpenStudy (lilsis76):

i dont have an equatioN haha says determine all of the solutions in the interval 0< Q<360 cos(Q/2) = 1/2 so the answers would be pi/3 and 2pi/3? right?

OpenStudy (lilsis76):

@anikay

OpenStudy (lilsis76):

@zepdrix is it pi/3 and 2pi/3?

zepdrix (zepdrix):

If the problem had been,\[\large \cos\left(\theta\right)=\frac{1}{2}\]We would have concluded that,\[\large \theta=\quad 60^o,\quad 300^o\] But our problem involved Q/2, not Q. Giving us,\[\large \cos\left(\frac{\theta}{2}\right)=\frac{1}{2}\] \[\frac{\theta}{2}=\quad 60^o,\quad 300^o\] From here, if we multiply both sides by 2, we can solve for Q.

OpenStudy (anikay):

yup^

zepdrix (zepdrix):

I'm a little confused why you're messing with the radians :C They're a little more confusing to work with. The domain of Q was GIVEN to us in DEGREES. So let's work in degrees! :D

OpenStudy (lilsis76):

AH! so i was right?! the 60 and 300?!

zepdrix (zepdrix):

\[\large \theta=\quad 120^o, \quad 600^o\]

OpenStudy (anikay):

and since we're going only to 360?

OpenStudy (anikay):

(radians are easier imo but degrees work k)

OpenStudy (lilsis76):

AH!!! im so lost haha

zepdrix (zepdrix):

aw :D

OpenStudy (anikay):

its almost like you're my lilsis or somethin xD

OpenStudy (lilsis76):

lol,

OpenStudy (lilsis76):

okay so i get the 1/2 and that is in the first quad giving me 60 degrees right?

OpenStudy (lilsis76):

@anikay

OpenStudy (anikay):

yes, and the 300

OpenStudy (lilsis76):

okay, how can i figure out the 300?!

OpenStudy (anikay):

it's a reference angle of 60

OpenStudy (lilsis76):

yes it is a ref angle 60

OpenStudy (lilsis76):

WAIT!!! 300 cuz its cos 1.2, sin - sqrt3 /2 ?!

OpenStudy (lilsis76):

right?! dang i think im slow

OpenStudy (anikay):

uh

OpenStudy (lilsis76):

haha did i do it right?!

OpenStudy (anikay):

yes

OpenStudy (lilsis76):

YAY!!!! > . < THANK YOU SO MUCH!!! ugh....

OpenStudy (anikay):

however, you have to double those and since you only go to 360, 120 is your only answer

OpenStudy (phi):

sometimes sketching the curve is helpful